# X-TIME: An in-memory engine for accelerating machine learning on tabular data with CAMs

### v3

@@ Abstract
Structured, or tabular, data is the most common format in data science. While deep learning models have proven formidable in learning from unstructured data such as images or speech, they are less accurate than simpler approaches when learning from tabular data. In contrast, modern tree-based Machine Learning (ML) models shine in extracting relevant information from structured data. An essential requirement in data science is to reduce model inference latency in cases where, for example, models are used in a closed loop with simulation to accelerate scientific discovery. However, the hardware acceleration community has mostly focused on deep neural networks and largely ignored other forms of machine learning. Previous work has described the use of an analog content addressable memory (CAM) component for efficiently mapping random forests. In this work, we focus on an overall analog-digital architecture implementing a novel increased precision analog CAM and a programmable network on chip allowing the inference of state-of-the-art tree-based ML models, such as XGBoost, CatBoost, and others. Results evaluated in a single chip at 16nm technology show 119$\times$ higher throughput at 9740$\times$ lower latency compared with a state-of-the-art GPU, with a 19W peak power consumption.
Index Terms:
@@
In-memory computing, tree-based ML, content addressable memories
I Introduction
@@
Extracting relevant information from structured (i.e., tabular) data is paramount in practical data science. Tabular data consist of a matrix of samples organized in rows, each of which has the same set of features in its columns, and is commonly used for medical, financial, scientific, and networking applications, enabling efficient search operations on large-scale datasets. While deep learning has shown impressive improvements in accuracy for unstructured data, such as image recognition or machine translation, tree-based machine learning (ML) models still outperform neural networks in classification and regression tasks on tabular data [1]. This difference in model accuracy is due to robustness to uninformative features of tree-based models, bias towards a smooth solution of neural networks models, and in general the need for rotation invariant models where the data is rotation invariant itself, as in the case of tabular data [1]. With high model accuracy, ease of use, and explainability [2] tree-based ML models are greatly preferred by the data science community [3], with $>74\%$ of data scientists preferring to use these models while $<40\%$ choose neural networks in a recent survey.
@@
However, tree-based models have garnered little attention in the architecture and custom accelerator research fields, particularly compared to deep learning. While small models are easily executed, performance issues for larger tree-based ML models arise due to limited inference latency both on CPU and GPUs [4, 5] originating from irregular memory accesses and thread synchronization issues. As large tree-based models are finding performance limits in existing hardware, larger models ($>$1M nodes) are demonstrating increasingly compelling uses in practical environments where high throughput and moderate latency are desired. For example, the winner of a recent IEEE-CIS Fraud detection competition used a tree-based ML model with 20M nodes [6]. Moreover, reducing inference latency is critical in scientific applications where for use in real-time processing, model latency must be $\sim$100 ns [7]. Similar low latency requirements for model inference arise in a real-time in-the-loop decision, data filtering systems, or streaming data scenarios, which are common across financial, medical, and cybersecurity domains [8, 6] where tabular data and tree-based ML also dominate. Thus, these models are reaching sizes and complexity levels for meaningful real-world tasks, as they are hitting limits on inference latency and throughput in current systems. Recently, custom accelerators for tree-based ML inference have been developed based either on FPGA [7, 8] or custom ASICs [9, 5]. A common architecture in most of the accelerators mentioned above implements look-up tables (LUTs) storing the attributes of a decision tree (DT) accessed multiple times during inference, with reads based on decision paths taken during the tree traversal. Along with low latency inference, an efficient hardware representation of trees is necessary to implement the required very large models.
@@
One promising new architectural approach for improving tree-based model inference is in-memory computing, where data are stored and processed in the same location. This approach eliminates access, transfer, and processing of data in a separate compute unit [10], reducing latency. Previous research has shown that these devices can efficiently perform multiplication functions in crossbar architectures [11]. However, crossbar architectures are limited to matrix operations, which are not useful for inference on tree-based models. Another class of in-memory computing primitives is Content Addressable Memories (CAMs) [12, 13], high-speed and highly parallel circuits that search for a given input in a table of stored words and return the matched search input’s location. More recently, analog CAMs have been proposed [14] which leverage the analog states in memristor devices. An analog CAM cell performs in situ comparisons of stored (analog) ranges with (analog) inputs, by computing if the input is within the range. Effectively, analog CAM arrays perform a massively parallel comparison between an input vector of analog values and stored words of analog ranges. Given that DTs are a set of conditional branches, analog CAMs naturally map DTs and perform highly parallel inference of decision forests [15]. Previous work showed the potential of this approach but was limited to component-level estimation, 4-bit precision, and random forest models [15].
@@
This work presents X-TIME, an in-memory engine for accelerating tree-based ML models with analog CAMs. In this work we demonstrate the need for a holistic architecture design, 8-bit operation, and support for multiple tree-based ML models, overcoming the limitation of previous work [15]. The major contributions of this paper compared to previous work [15] are (1) a novel analog CAM cell design and operation to perform decision tree inference at doubled (i.e. 8-bit) memristor precision (i.e. 4-bit), (2) an in-core and in-network leave accumulation architecture to seamlessly perform inference of a large number of trees in parallel, (3) a programmable Network-on-Chip (NoC) to operate any tree-based ML model regardless of the ensemble reduction operation, (4) a compiler supporting the mapping of multiple tree-based ML models and (5) a rigorous characterization method and scaling comparison with respect to GPUs, FPGAs, and conventional digital accelerators. Our simulated evaluation shows up to 119$\times$ increased throughput at 9740$\times$ decreased latency in a single 19W chip.
II Background
@@
II-A Tree-based Machine Learning
@@
II-B Tree-based ML model inference on GPU
While training DT models is relatively fast when compared to, for example, a large deep neural network, the computations involved in inference are very irregular and not particularly suitable for CPUs and GPUs, becoming inefficient for very large models, with $>1$ M nodes [6]. While CPUs can run irregular control paths fairly well, their parallelism is limited. GPUs have a very large parallelism, but suffer from the irregularity of memory access patterns, as will be explained in detail below. Recent NVIDIA libraries implementations have shown order of magnitude improvements compared to CPU and previous GPU implementations  [19]. However, three different factors still limit GPU throughput and latency [4], in particular for tree ensembles. First, a high number of uncoalesced memory accesses, due to the low ratio between requested and fetched data, causes low memory efficiency. Trees are usually stored in memory with nodes from different trees interleaved in breadth-first order, resulting in coalesced memory accesses close to the root, where children nodes are more likely to be part of the same memory transaction of the parent node, but uncoalesced memory accesses further from the root. Thus, with increasing tree depth the proportion of coalesced memory accesses decreases and memory system efficiency also decreases. Second, synchronization between threads and the consequent load imbalance due to waiting for the deepest trees limits latency. A common implementation scheme on GPU assigns each tree in the ensemble to a different thread by round-robin. At the end of each thread block, a reduction operation is then performed. However, a given ensemble may contain short and tall trees, with the latter having substantially larger latency. Thus, each thread block must wait until the slowest (tallest) tree has finished processing, increasing overall latency. Finally, with a larger number of trees $N_{trees}$ in the ensemble, the increasing number of thread blocks and the global reduction performed between thread blocks becomes a significant overhead. Multiple techniques have been proposed to increase the throughput of tree ensemble inference [19, 4], mainly by re-organizing the memory - for example, for sparse trees and scheduling of threads - but with limited impact on latency. In this work, we focus on addressing these limitations with an in-memory computing architecture using content-addressable memories.
@@ II-B Tree-based ML model inference on GPU
II-C Analog Content Addressable Memory
CAMs are specialized memory devices that search a table of stored words (rows) or keys in parallel for a given input query. The CAM search output is a one-hot vector where a high value corresponds to a match between the input data query and stored row word key [12, 13]. In conventional CAMs, this output vector is converted into a single address by a priority encoder. Each CAM cell functions as both memory and bit comparison, and the input query is compared to each CAM word simultaneously to produce a high throughput, low latency few-cycle compare operation at the cost of high power and area. In addition to binary (0 and 1) values, ternary CAMs (TCAMs) have a third wildcard value, X (or “don’t care”), which can be used to compress the stored CAM table or perform partial match searches. When an X is stored or searched, that cell matches irrespective of whether the searched or stored value is 1 or 0. Fig. 2(a) shows a schematic illustration of a conventional CAM/TCAM circuit, where input queries are applied on the column search lines (SLs) and outputs are read along the row ”match lines” (MALs). An XNOR operation occurs between the query values and stored cell values in the corresponding column. The MAL $j_{th}$ row output is produced by a bit-wise AND operation between cells along the row and returns a match (high) or mismatch (low). The red boxes in Fig. 2(a) correspond to mismatches between input and stored value. In practice, MALs are initially pre-charged high and a mismatch in a cell activates a pull-down transistor connected to MAL to discharge it, such that a match is returned only if all pull-down transistors remain off during a search operation.
TCAMs are used ubiquitously in networking applications that require high-throughput and low-latency operations, for example for packet classification and forwarding, and access control lists [12]. While offering high performance, conventional SRAM-TCAMs are power and area-hungry and require 16 transistors. Recently, there has been significant interest in developing TCAMs with memristor devices and resistive non-volatile memories [20, 21, 22] with compact size and low-power operation. As programming these devices is relatively slow, emerging TCAMs have targeted in-memory computing applications [23, 13] where stored value updates are infrequent.
Beyond direct digital memory operation, memristor devices have also demonstrated ranges of stable analog states of $>16$ levels. Leveraging this analog state capability, recent work proposed and demonstrated an analog CAM circuit design and operation [14] (conceptual representation in Fig 2(b)). In an analog CAM, ranges are stored in each cell, an analog input is applied vertically along columns and a match is returned along rows if the inputs are all within the stored range of each cell in a row. An analog CAM row implements
@@ II-C Analog Content Addressable Memory
where $q$ is the query input vector applied to DL, and $T_{L}$ and $T_{H}$ are the stored range thresholds for the CAM cell range lower and upper bounds, respectively. In the analog CAM circuit, no analog-to-digital conversion is required, eliminating an expensive component which is a limiting bottleneck of other analog in-memory computing primitives [24]. Multiple analog CAM circuits have been realized using different technologies such as memristor devices [14] and ferroelectric memory devices [25]. In such circuits, two analog memory elements represent the lower and upper bound thresholds. Analog CAMs have been shown to map CAM tables more efficiently [14], as well as implement novel functions, such as computing distances for memory-augmented neural networks [25, 22], decision tree inference [15] and pattern learning [26].
@@
II-D Mapping tree models to analog CAMs
@@
The X-TIME compiler takes as input the description of a DT ensemble in a tabular mode, where each row contains the node attributes (feature ID and value) and the left/right children. The compiler traverses the tree structures extracting all the root-to-leaf paths. Finally, it maps each of them to an ACAM row. As an example, Fig. 3 shows the mapping of Fig. 1(a). Each column corresponds to a different feature of the input vector $q$. Each node is mapped in the column corresponding to the feature used in the node’s comparison. The comparison’s threshold value is programmed in the analog CAM’s lower and/or upper bound, according to the direction, left or right, taken by the comparison’s result for the root-to-leaf path. A “don’t care” is used if a feature is missing, by programming the full range (0 to 1 in our example) in the analog CAM cell. The feature vector (query) is applied on the DL and the output on the MAL is high for the matched branch, i.e. the traversed root-to-leaf path, in a single CAM operation. This mapping requires the number of CAM rows to match the number of tree leaves, and the number of CAM columns to match the number of features $N_{feat}$. MALs can be connected directly to the Word Lines (WLs) of a conventional RAM storing leaf values to return the model’s prediction. Effectively, the analog CAM traverses root-to-leaf paths, and the RAM retrieves leaf values.

### v4

@@ Abstract
Structured, or tabular, data is the most common format in data science. While deep learning models have proven formidable in learning from unstructured data such as images or speech, they are less accurate than simpler approaches when learning from tabular data. In contrast, modern tree-based Machine Learning (ML) models shine in extracting relevant information from structured data. An essential requirement in data science is to reduce model inference latency in cases where, for example, models are used in a closed loop with simulation to accelerate scientific discovery. However, the hardware acceleration community has mostly focused on deep neural networks and largely ignored other forms of machine learning. Previous work has described the use of an analog content addressable memory (CAM) component for efficiently mapping random forests. In this work, we develop an analog-digital architecture that implements a novel increased precision analog CAM and a programmable chip for inference of state-of-the-art tree-based ML models, such as XGBoost, CatBoost, and others. Thanks to hardware-aware training, X-TIME reaches state-of-the-art accuracy and 119$\times$ higher throughput at 9740$\times$ lower latency with >150×>150$\times$> 150 × improved energy efficiency compared with a state-of-the-art GPU for models with up to 4096 trees and depth of 8, with a 19W peak power consumption.
{IEEEkeywords}
@@
Content Addressable Memories, Decision Trees, In-memory Computing, ReRAM
1 INTRODUCTION
[…]
Extracting relevant information from structured (i.e., tabular) data is the foundation of practical data science. Tabular data consists of a matrix of samples organized in rows, each of which has the same set of features in its columns, and is commonly used for medical, financial, scientific, and networking applications, enabling efficient search operations on large-scale datasets, making them one of the most common sources of data in the data science industry. Moreover, data collected in real-time for heterogenous sources, for example during scientific experiments, are organized in tabular data and with fast processing may lead to in-the-loop intelligent control.
@@ 1 INTRODUCTION
While deep learning has shown impressive improvements in accuracy for unstructured data, such as image recognition or machine translation, tree-based machine learning (ML) models still outperform neural networks in classification and regression tasks on tabular data [1], including modern Transformers for tabular data [2]. This difference in model accuracy is due to robustness to uninformative features of tree-based models, bias towards a smooth solution of neural networks models, and in general the need for rotation variant models where the data is rotation variant itself, as in the case of tabular data [1]. With high model accuracy, ease of use, and explainability [3] tree-based ML models are greatly preferred by the data science community [4], with $>74\%$ of data scientists preferring to use these models while $<40\%$ choose neural networks in a recent survey.
@@
However, tree-based models have garnered little attention in the architecture and custom accelerator research fields, particularly compared to deep learning. While small models are easily executed, performance issues for larger tree-based ML models arise due to limited inference latency both on CPU and GPUs [5, 6] originating from irregular memory accesses and thread synchronization issues. As large tree-based models ($>$1M nodes) are finding performance limits in existing hardware, they are demonstrating increasingly compelling uses in practical environments where high throughput and moderate latency are desired. For example, the winner of a recent IEEE-CIS Fraud detection competition used a tree-based ML model with 20M nodes [7]. Moreover, reducing inference latency is critical in scientific applications where for use in real-time processing, model latency must be $\sim$100 ns [8]. Similar low latency requirements for model inference arise in a real-time in-the-loop decision, data filtering systems, or streaming data scenarios, which are common across financial, medical, and cybersecurity domains [9, 7] where tabular data and tree-based ML also dominate. Thus, these models are reaching sizes and complexity levels for meaningful real-world tasks, as they are hitting limits on inference latency and throughput in current systems. Recently, custom accelerators for tree-based ML inference have been developed based either on FPGA [8, 9] or custom ASICs [10, 6]. A common architecture in most of the accelerators mentioned above implements look-up tables (LUTs) storing the attributes of a decision tree (DT) accessed multiple times during inference, with reads based on decision paths taken during the tree traversal. Along with low latency inference, an efficient hardware representation of trees is necessary to implement the required very large models.
@@
One promising new architectural approach for improving tree-based model inference is in-memory computing, where data are stored and processed in the same location. This approach eliminates access, transfer, and processing of data in a separate compute unit [11], reducing latency. Previous research has shown that multiplication functions can efficiently be performed in crossbar architectures [12]. However, crossbar architectures are limited to matrix operations, which are not useful for inference on tree-based models. Another class of in-memory computing primitives is Content Addressable Memories (CAMs) [13, 14], high-speed and highly parallel circuits that search for a given input in a table of stored words and return the matched search input’s location. CAMs use relatively cheap peripherals, e.g. sense amplifier circuits rather than expensive ADCs required by crossbar arrays for in-memory computing operations. More recently, analog CAMs have been proposed [15] which leverage the analog states in resistive switching memories (RRAM) [16]. An analog CAM performs the comparisons of stored (analog) ranges with (analog) inputs, by computing if the input is within the range. Given that DTs are a set of conditional branches, analog CAMs naturally map DTs and perform highly parallel inference of decision forests [17, 18]. Our previous work showed the potential of this approach but was limited to component-level estimation, 4-bit precision, and random forest models [17], while the full architecture study in this work - which includes designs for higher precision Analog CAM operation, study of peripheral circuits and builds in support for multiple ML models convincingly demonstrates the feasibility of the initial concept.
@@
This work presents X-TIME, an in-memory engine for accelerating tree-based ML models with analog CAMs. In this work we demonstrate the need for a vertical architecture design, 8-bit operation, and support for multiple tree-based ML models, overcoming the limitations of previous work [17]. The major contributions of this paper compared to previous work [17] are at the circuit level(1) a novel analog CAM cell design and operation to perform decision tree inference at doubled (i.e. 8-bit) resistive memory precision (i.e. 4-bit), (2) a complete design of the core and network operation, and at the architecture level (3) an end-to-end design targeting a large scale accelerator, (4) a compiler (X-TIME-C) supporting the mapping of multiple tree-based ML models to the HW starting from established intermediate representation [19], and at the co-design level (5) an HW-aware model training algorithm for reaching state of the art accuracy and (6) a complete simulation framework to benchmark and rigorous characterization comparison against NVIDIA V100 GPUs, FPGAs, and conventional digital accelerators. Our simulated evaluation shows up to 119$\times$ increased throughput at 9740$\times$ decreased latency with >150×>150$\times$> 150 × improved energy efficiency in a single 19W chip.
2 BACKGROUND
@@
2.1 Tree-based Machine Learning
@@
While training DT models is relatively fast when compared to, for example, a large deep neural network, the computations involved in inference are very irregular and not particularly suitable for CPUs and GPUs [6], becoming inefficient for very large models, with $>1$ M nodes [7]. While CPUs can run irregular control paths fairly well, their parallelism is limited. GPUs have a very large parallelism, but suffer from the irregularity of memory access patterns, as explained in detail in Supplementary Information LABEL:supp-si:section:gpu_inference.
@@
2.2 Analog Content Addressable Memory
CAMs are specialized memory devices that search a table of stored words (rows) or keys in parallel for a given input query. CAMs output a one-hot vector where a high value corresponds to a match between the input data query and stored row word key [13, 14]. In conventional CAMs, this output vector is converted into a single address by a priority encoder. Each CAM cell functions as both memory and bit comparison, and the input query is compared to each CAM word simultaneously to produce a high throughput, low latency few-cycle compare operation at the cost of high power and area. In addition to binary (0 and 1) values, ternary CAMs (TCAMs) have a third wildcard value, X (or don’t care), which can be used to compress the stored CAM table or perform partial match searches. When an X is stored or searched, that cell matches irrespective of whether the searched or stored value is 1 or 0. Fig. 2(a) shows a schematic illustration of a conventional CAM/TCAM circuit, where input queries are applied on the column Search Lines (SLs) and outputs are read along the row Match Lines (MALs). Essentially, an XNOR operation occurs between the query values and stored cell values in the corresponding column. The MAL $j_{th}$ row output is produced by a bit-wise AND operation between cells along the row and returns a match (high) or mismatch (low). The red boxes in Fig. 2(a) correspond to mismatches between input and stored value. In practice, MALs are initially pre-charged high and a mismatch in a cell activates a pull-down transistor connected to MAL to discharge it, such that a match is returned only if all pull-down transistors remain off during a search operation.
TCAMs are used ubiquitously in networking applications that require high-throughput and low-latency operations, for example for packet classification and forwarding, and access control lists [13]. While offering high performance, conventional SRAM-TCAMs are power and area-hungry and require 16 transistors. Recently, there has been significant interest in developing TCAMs with RRAM devices  [23, 24, 25] with compact size and low-power operation. As programming these devices is relatively slow, emerging TCAMs have targeted in-memory computing applications [26, 14] where stored value updates are infrequent.
Beyond direct digital memory operation, RRAM devices have also demonstrated ranges of stable analog states of $>16$ levels [27]. Leveraging this analog state capability, recent work proposed and demonstrated an analog CAM circuit design and operation [15] (conceptual representation in Fig 2(b)). In an analog CAM, ranges are stored in each cell, an analog input is applied vertically along columns and a match is returned along rows if the inputs are all within the stored range of each cell in a row. An analog CAM row implements
@@ 2.2 Analog Content Addressable Memory
where $q$ is the query input vector applied to Data Line (DL), and $T_{L}$ and $T_{H}$ are the stored range thresholds for the CAM cell range lower and upper bounds, respectively. In the analog CAM circuit, no analog-to-digital conversion is required, eliminating an expensive component which is a limiting bottleneck of other analog in-memory computing primitives [28]. Multiple analog CAM circuits have been realized using different technologies such as RRAM devices [15] and ferroelectric memory devices [29]. In such circuits, two analog memory elements represent the lower and upper bound thresholds. Analog CAMs have been shown to map CAM tables more efficiently [15], as well as implement novel functions, such as computing distances for memory-augmented neural networks [29, 25], decision tree inference [17] and pattern learning [30].
@@
2.3 Mapping tree models to analog CAMs
@@
Fig. 3 shows the mapping of the decision tree in Fig. 1(a) to the Analog CAM. Each row represents a branch of the tree and each column corresponds to a different feature of the input vector $q$. Each node is mapped in the column corresponding to the feature used in the node’s comparison. The comparison’s threshold value is programmed in the analog CAM’s lower and/or upper bound, according to the direction, left or right, taken by the comparison’s result for the root-to-leaf path. A don’t care symbol is used if a feature is missing, by programming the full range (0 to 1 in our example) in the analog CAM cell. The feature vector (query) is applied on the DL and the output on the MAL is high for the matched branch, i.e. the traversed root-to-leaf path, in a single CAM operation. This mapping requires the number of CAM rows to match the number of tree leaves, and the number of CAM columns to match the number of features $N_{feat}$. MALs can be connected directly to the Word Lines (WLs) of a conventional SRAM storing leaf values to return the model’s prediction. Effectively, the analog CAM traverses root-to-leaf paths, and the RAM retrieves leaf values.